OpenStudy (anonymous):

The position x, in meters, of an object moving along the x-axis is given as a function of time t, in seconds, as x(t)=0.0981t^4-0.231t^3+0.773t^2+5.83t-9.27. Find the object's acceleration at time t=3.45s

OpenStudy (abhisar):

\(\huge \sf Acceleration = \huge \frac{d(velocity)}{dt}\)

OpenStudy (abhisar):

So first we will have to calculate the velocity as we calculated in the last question. Can you calculate the equation for velocity ?

OpenStudy (abhisar):

\(\sf \huge Velocity = \huge \frac{dx}{dt}\)

OpenStudy (anonymous):

would the 9.27 be negative? when subtracting it from .585?

OpenStudy (abhisar):

\(\sf = 0.3924t^3-0.693t^2+1.546t\) This is velocity

OpenStudy (abhisar):

Now u'll have to differentiate it once more to get an equation for acceleration. Can you ?

OpenStudy (anonymous):

ok so,-0.231t^3+0.773t^2+583t = the velocity?

OpenStudy (abhisar):

Wr did you got this from ?

OpenStudy (anonymous):

from the problem

OpenStudy (anonymous):

or is x the total velocity?

OpenStudy (abhisar):

x is position/Displacement

OpenStudy (abhisar):

You differentiate it to get velocity.

OpenStudy (abhisar):

Then differentiate velocity to get acceleration.

OpenStudy (abhisar):

You know how to differentiate ?

OpenStudy (anonymous):

not exactly

OpenStudy (abhisar):

Haven't u learnt about calculus ?

OpenStudy (anonymous):

i've just started calculus, we are currently in limits

OpenStudy (abhisar):

ok, i am simplifying it for you though u'll learn it in detail

OpenStudy (anonymous):

ok,thanks

OpenStudy (abhisar):

If i say you to differentiate \(\sf 5t^n\) with respect to t, then the answer will be \(\sf 5nt^{n-1}\)

OpenStudy (abhisar):

Example: Differentiate \(\sf 5t^7\) with respect to t Ans: \(\sf 35t^6\)

OpenStudy (abhisar):

Got it ?

OpenStudy (anonymous):

so you multiplied the 7 to the 5 and subtracted 1 from the 7?

OpenStudy (abhisar):

yes

OpenStudy (abhisar):

Now differentiate \(\sf 8t^9\) wrt t

OpenStudy (anonymous):

72t^8

OpenStudy (abhisar):

\(\huge\checkmark\)

OpenStudy (abhisar):

Differentiation of a constant term is always zero

OpenStudy (abhisar):

Example if i say you to differentiate 36 wrt t, then the answer will be 0. Got it ?

OpenStudy (anonymous):

ok, yes...if the number is negative, for example -4t^3, would the 4 remain negative or should it be positive?

OpenStudy (abhisar):

It will remain negative

OpenStudy (abhisar):

Now, Differentiate \(\sf 3t^7-5t^4+68\) wrt t

OpenStudy (anonymous):

21t^6-20t^3+68

OpenStudy (abhisar):

68 will become 0 after differentiation. It's a constant term

OpenStudy (anonymous):

oh ok, i see

OpenStudy (abhisar):

One more time, Differentiate \(\sf 45x^4+78x^3-8.9t+56\) wrt x

OpenStudy (anonymous):

180t^3+234t^2-8.9t

OpenStudy (abhisar):

\(\sf 8.9x\) can be written as \(\sf 8.9x^1\). So when you will differentiate it wrt to x it will become 8.9

OpenStudy (anonymous):

ok, i was unsure about that

OpenStudy (abhisar):

also you can not change x with t

OpenStudy (anonymous):

oh sorry i didn't realize i did that...

OpenStudy (abhisar):

Answer will be \(\sf 180x^3+234x^2-8.9~and~not~180t^3+234t^2-8.9t\)

OpenStudy (abhisar):

are you sure you can do the next one correctly in one attempt ?

OpenStudy (anonymous):

yeah

OpenStudy (abhisar):

allryt let's see Differentiate \(\sf 0.0981t^4-0.231t^3+0.773t^2+5.83t-9.27\)

OpenStudy (anonymous):

0.3924t^3-0.693t^2+1.554t^1+5.83t

OpenStudy (abhisar):

\(\sf 0.3924t^3-0.693t^2+1.546t^1+5.83\)

OpenStudy (anonymous):

how does the t on 5.83 cancel out?

OpenStudy (abhisar):

Differentiation of 5.83t = ?

OpenStudy (anonymous):

5.83

OpenStudy (abhisar):

yes, 5.83t = \(\sf 5.83t^1\), when u differentiate it, \(\sf 1\times 5.83t^{1-1}=5.83t^0=5.838\)

OpenStudy (anonymous):

so from .3924t^{3}-.693t ^{2}+1.544t+5.83 you plug in the time 3.45s for t? .....oh yes ok i see, i'm sorry my mistake

OpenStudy (abhisar):

No, you need to differentiate it once more

OpenStudy (abhisar):

Differentiate \(\sf 0.3924t^3−0.693t^2+1.546t^1+5.83\)

OpenStudy (anonymous):

1.1772t^2-1.386t+1.544

OpenStudy (anonymous):

how do you know when you are done differentiating? is it when it gets into slop-point form?

OpenStudy (anonymous):

or is it when t is simplified?

OpenStudy (abhisar):

When you differentiate displacement you get velocity. When you differentiate velocity you get acceleration. So if you have to calculate acceleration from displacement you need to differentiate it twice. Got it ?

OpenStudy (abhisar):

First differentiation will give you velocity and the second one will give acceleration. Ryt ?

OpenStudy (anonymous):

oh ok...

OpenStudy (anonymous):

so it should be 1.1772t^2-1.386t+1.544 and plug in 3.45 for t right?

OpenStudy (abhisar):

yes

OpenStudy (anonymous):

thank you so much!!!